OpenStudy (anonymous):

Complete the square to solve the equation below? Check all that apply. x^2-10x-5=9

OpenStudy (anonymous):

A.\[5+\sqrt{39}\]

OpenStudy (anonymous):

\[x^2-10x=14\]

OpenStudy (anonymous):

B.\[-10-\sqrt{24}\]

OpenStudy (anonymous):

C.\[5-\sqrt{39}\]

OpenStudy (anonymous):

D.\[10+\sqrt{24}\]

OpenStudy (anonymous):

so can someone help

OpenStudy (ybarrap):

$$x^2-10x-14=0\\ (x-10/2)^2-14-(10/2)^2=0\\ =(x-5)^2-39=0\\ =(x-5)^2=39\\ (x-5)=\pm\sqrt{39}\\ x=5\pm\sqrt{39}$$